OpenStudy (anonymous):

Solve the system using the substitution method: x+2y=6 3x-2y=2

OpenStudy (anonymous):

so the substitution method states that you need to subsitute one variable for another. Lets substitue for y, how do we do this? Solve for y with ethier equation.

OpenStudy (anonymous):

i don't know!

OpenStudy (anonymous):

I didnt ask you a questions lol I told u to solve for y on one of those equations I wont help if you dont try

OpenStudy (anonymous):

i've been sitting here for like half an hour looking through the steps on how to do it

OpenStudy (anonymous):

take one of those equations and solve for y, you have done this before.

OpenStudy (anonymous):

I did, but i ended up with -4y= -3x+2....and that's not right

OpenStudy (anonymous):

it would be -2y=-3x+2 where did u get the 4?

OpenStudy (anonymous):

also thats not finished you need to get the 2 away from the y

OpenStudy (anonymous):

i have know idea and ok

OpenStudy (anonymous):

3x-2y=2 you subtract 3x -2y=-3x+2 then what?

OpenStudy (anonymous):

3x-2y=2?

OpenStudy (anonymous):

that is the original equation you divide both sides by 2 to get y by itself just like we did on that other paper

OpenStudy (anonymous):

3x+y=1

OpenStudy (anonymous):

i meant -1

OpenStudy (anonymous):

0.o no were at -2y=-3x+2 if you divide -2 you get y=3/2x+1

OpenStudy (anonymous):

that's what i had at first, but then you said the original equation was something different

OpenStudy (anonymous):

I did not and you never had that you had 3x+y=1 you still need to solve for y

OpenStudy (anonymous):

now that you have solved for y uou can plug that equation in for the y in your second equation

OpenStudy (anonymous):

x+2y=6 3x-2y=2 im gonna use the top one becuase its easier but x+2y=6 -x -x 2y=-x+6 /2 /2 y=(-1/2)x+3 now take the other equation solve for y and SUBSTITUTE the Y;s y=(-1/2)x+3 then solve for 3x-2y=2 -2y=-3x+2 /-2 /-2 y=(3/2)x+1 y=y (3/2)x+1 = (-1/2)x+3 (4/2)x=2 *2 *2 4x=4 x=1

OpenStudy (anonymous):

thanks